OpenStudy (anonymous):

solve this system of equation(cannot be a fraction) x+y+z=1 2x+5y+2z=5 -x+6y-3z=-2 please show steps

OpenStudy (anonymous):

r1:x+y+z=1 r2:2x+5y+2z=5 r3: -x+6y-3z=-2 (r2-2r1) 3y = 3 y=1 now substitute 1 for y in r1 and r3 r4:x+z=0 r5: -x -3z = -8 (r5+r4) -2z = -8 z=4 now substitute 4 for z in r4 x+4=0 x=-4